# MLC at Feigenbaum points

### v2

@@ 3.3.4. Canonical arc diagram
Recall that an arc $\alpha$ is a proper path in $V$ up to proper homotopy. Any arc $\alpha$ connects some components $J_{k}$ and $J_{i}$ of $\partial V$ and lifts to an arc $\alpha$ in ${\mathbb{D}}$ connecting some intervals ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$. We let ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})={\mathcal{W}}_{\overline{\mathbb{D}}}({\mathbf{J}}_{k},{\mathbf{J}}_{i})$ to be the width of all curves in $\mathbb{D}$ connecting ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$. Then ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})$ is also the width of the rectangle ${\boldsymbol{\Pi}}_{\mbox{\boldmath$\displaystyle\alpha$}}=\overline{\mathbb{D}}$ with horizontal sides ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$. It is independent of the lift used.
If ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})>2$, then by removing from ${\boldsymbol{\Pi}}_{\mbox{\boldmath$\displaystyle\alpha$}}$ the square buffers we obtain a rectangle ${\boldsymbol{\Pi}}^{{\operatorname{can}}}_{\mbox{\boldmath$\displaystyle\alpha$}}$. In this case we let
${\mathcal{W}}^{\operatorname{can}}(\alpha)={\mathcal{W}}^{\operatorname{can}}({{\mbox{\boldmath$\displaystyle\alpha$}}})={\mathcal{W}}({\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{\boldmath$\displaystyle\alpha$}})={\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})-2.$
By construction, the $\Delta$-orbit of ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{\boldmath$\displaystyle\alpha$}}$ consists of pairwise disjoint rectangles. Therefore, ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{\boldmath$\displaystyle\alpha$}}$ projects injectively onto a rectangle ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha}$ in $V$. Moreover, different ${\boldsymbol{\Pi}}^{\operatorname{can}}_{{\mbox{\boldmath$\displaystyle\alpha$}}_{i}}$ project to pairwise disjoint ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha_{i}}$.
If ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})\leq 2$ we let ${\mathcal{W}}^{\operatorname{can}}(\alpha)=0$ and ${\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)\coloneqq\emptyset$. Arcs with positive weight form the canonical weighted arc diagram (WAD) on $V$:
@@ 3.3.5. Local WADs and the thick-thin decomposition
See [L3, §7.6.3] for a reference of the following fact.
@@ 3.3.8. WAD under covering
Assume that $f\colon U\to V$ as a finite-degree covering between surfaces with boundaries. Then, see [K, Lemma 3.3]
@@ 3.4.2. Definitions
A ${}^{\bullet}$-quadratic-like map (“$\psi^{\bullet}$-ql map”) is a pair of holomorphic maps
@@
    $f\colon U\to V$ is a double branched covering (we usually normalize it so that its critical point is located at 00 0);
@@ 3.4.3. $\psi^{\bullet}$-ql renormalization
$U_{n,i}\coloneqq{\mathbb{A}}\left(U\setminus f^{-1}({\boldsymbol{\mathfrak{B}}}^{[n]}),\ \partial{\mathfrak{B}}^{[n](1)}_{i}\right)\bigcup_{\partial{\mathfrak{B}}^{[n](1)}_{i}}{\mathfrak{B}}^{[n](1)}_{i}$
@@
We will write suppress index “00 0” for $F_{n,0}$:
@@
[…]
For example, one can take $\Upsilon={\mathfrak{T}}$ or $\Upsilon={\mathfrak{T}}\cup{\boldsymbol{\mathfrak{B}}}^{[m]}$. To define “$\psi^{\Upsilon}$-renormalization”, one should extend a ql prerenormalization $f_{n,i}\colon{X_{n,i}\to Y_{n,i}}$ along all curves in $V\setminus\bigcup_{j}\Upsilon_{f_{n,j}}$. Moreover, $\Upsilon_{f_{n,j}}$ can often be enlarged, see §3.4.5.
@@ 3.4.4. Sup-Chain Rule for renormalization domains
    ${\boldsymbol{\Upsilon}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[n_{1}+n_{2}]}\cup\left({\boldsymbol{\mathfrak{B}}}^{[n_{2}]}\setminus{\mathfrak{B}}^{[n_{2}]}_{0}\right),$
@@ 3.4.5. Unbalanced renormalization from $\psi^{\bullet}$ to $\psi$-ql maps
We recall from §2.4.5 that the set ${\boldsymbol{\Upsilon}}$ is inhomogeneous: $f_{X}$ does not permute components of ${\boldsymbol{\Upsilon}}$
@@ 3.4.7. Sup-Chain rule for iterations
[…]
@@ 3.5.2. WADs of $F$
Assume that ${\boldsymbol{\mathfrak{B}}}^{[n]}$ is well-defined, i.e., $F$ is $n+1$ renormalizable. Let us consider the canonical WAD of $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],m}$:
${\mathcal{A}}^{(m),k}\equiv{\mathcal{A}}^{[n],(m),k}\coloneqq{\operatorname{WAD}}\left(U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],m}\right).$
@@
  • ${\mathcal{A}}_{i}^{(m),k}$ to be the local WAD for ${\mathfrak{B}}^{[n],(m),k}_{i}$, see §3.3.5: ${\mathcal{A}}_{i}^{k,(m)}$ consisting of arcs adjacent to ${\mathfrak{B}}^{[n],(m),k}_{i}$ such that the weights of self-arcs adjacent to ${\mathfrak{B}}^{[n],(m),k}_{i}$ are doubled;
@@
Assume that in the dynamical plane of $F={\mathcal{R}}^{n_{0}\bullet}\colon U\rightrightarrows V$, there is a non-trivial horizontal lamination of curves ${\mathcal{L}}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ emerging from ${\mathfrak{B}}^{[n]}_{0}$ and landing at ${\boldsymbol{\mathfrak{B}}}^{[n]}$. Then
@@ 3.5.3. WADs and ${\mathcal{W}}_{\bullet}(F_{n,i})$
(3.24) $\frac{1}{2}{\mathcal{W}}(F_{n,0})\leq{\mathcal{W}}(F_{n,i})\leq{\mathcal{W}}(F_{n,0})\hskip 19.91684pt\text{ for all }\ i.$
Indeed, let $\gamma_{i}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ be the peripheral hyperbolic geodesic around ${\mathfrak{K}}^{[n]}_{i}$. Then the hyperoblic length of $|\gamma_{i}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$ is proportional to ${\mathcal{W}}(F_{n,i})$. Let $\gamma^{1}_{i}$ be the lift of $\gamma_{i}$ under $f\colon U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}\to V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ so that $\gamma^{m}_{i}\subset U^{(m)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ is the peripheral hyperbolic geodesic around ${\mathfrak{B}}^{[n],(1)}_{i-1}$, where the subscript is $\operatorname{mod}p_{n}$. Counting the degree of $f\colon\gamma^{1}_{i}\to\gamma_{i}$, we obtain
  • $|\gamma^{1}_{i}|_{U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}=|\gamma_{i}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$ if $i>1$; and
  • $\frac{1}{2}|\gamma^{1}_{1}|_{U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}=|\gamma_{1}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$.
Finally, $|\gamma^{1}_{i}|_{U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}\geq|\gamma_{i-1}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$; this implies (3.24).
@@
Assume that $m\leq k$. By Lemma 3.6, the set ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\subset{\mathfrak{B}}_{F}$ is $\iota$-proper. Let $\gamma$ be a proper curve in $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ emerging from ${\mathfrak{B}}^{[n],(m)}_{a}$. Applying $\iota^{-s}$ along $\gamma$, we construct, see (3.26) below, its restriction $(\iota^{*})^{s}[\gamma]$ which is a proper curve in $U^{k+s}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$. Moreover,
@@
The curve $\gamma\in{\mathcal{F}}^{\operatorname{full}}(R)$ becomes vertical under the restriction by $(\iota^{s})^{*}$ if and only if $\gamma$ hits $\iota^{s}\big{(}\partial U^{k+s}\big{)}$. This is impossible because $\gamma_{1},\gamma_{2}$ do not hit $\iota^{s}\big{(}\partial U^{k+s}\big{)}$. ∎
@@ 3.5.6. Monotonicity of the ${\mathcal{A}}^{k}_{\mathrm{hor}}$
It follows essentially from Lemma 3.3 (by lifting $\iota$ to the universal coverings, see [K, §3.5] for reference) that
(3.27) ${\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}}\geq{\mathcal{A}}^{[n],(m),k+1}_{\mathrm{hor}}\hskip 14.22636pt\text{ and }\hskip 14.22636pt{\mathcal{A}}^{[n],(m),k}_{i,{\mathrm{hor}}}\geq{\mathcal{A}}_{i+1,{\mathrm{hor}}}^{[n],(m),k+1}.$
@@
It follows from Decomposition (3.28) that there is a constant $C\equiv C_{p_{n},m}$ such that for all $M>0$
[…]
in the following sense, see §3.3.6 for notation. For every arc $\alpha$ in ${\mathcal{A}}^{[n],(m+1),k}_{\mathrm{hor}}-{\operatorname{max}}\{M,1\}C$, there is a vertical sublamination
${\mathcal{L}}_{\alpha}\subset{\mathcal{F}}(e)\equiv{\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha){\ \ }{\ \ }\text{ with }{\ \ }{\ \ }{\mathcal{W}}({\mathcal{L}}_{\alpha})\geq{\mathcal{W}}(\alpha)-{\operatorname{max}}\{M,1\}C$
@@
(3.29) $\gamma_{i}\in{\mathcal{F}}^{\operatorname{full}}\left({\mathcal{A}}^{[n],(m+1),k}_{\mathrm{hor}}-C\right).$
The lamination ${\mathcal{L}}_{a}$ is constructed by removing all $\gamma$ from ${\mathcal{F}}(\alpha)$ that do not satisfy (3.29) – the width of removed curves is bounded by ${\operatorname{max}}\{M,1\}C$.

### v3

@@ 3.3.4. Canonical arc diagram
Recall that an arc $\alpha$ is a proper path in $V$ up to proper homotopy. Any arc $\alpha$ connects some components $J_{k}$ and $J_{i}$ of $\partial V$ and lifts to an arc $\alpha$ in ${\mathbb{D}}$ connecting some intervals ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$. We let ${\mathcal{W}}({{\mbox{$\displaystyle\alpha$}}})={\mathcal{W}}_{\overline{\mathbb{D}}}({\mathbf{J}}_{k},{\mathbf{J}}_{i})$ to be the width of all curves in $\mathbb{D}$ connecting ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$. Then ${\mathcal{W}}({{\mbox{$\displaystyle\alpha$}}})$ is also the width of the rectangle ${\boldsymbol{\Pi}}_{\mbox{$\displaystyle\alpha$}}=\overline{{\mathbb{D}}}$ with horizontal sides ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$. It is independent of the lift used.
If ${\mathcal{W}}({{\mbox{$\displaystyle\alpha$}}})>2$, then by removing from ${\boldsymbol{\Pi}}_{\mbox{$\displaystyle\alpha$}}$ the square buffers we obtain a rectangle ${\boldsymbol{\Pi}}^{{\operatorname{can}}}_{\mbox{$\displaystyle\alpha$}}$. In this case we let
${\mathcal{W}}^{\operatorname{can}}(\alpha)={\mathcal{W}}^{\operatorname{can}}({{\mbox{$\displaystyle\alpha$}}})={\mathcal{W}}({\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{$\displaystyle\alpha$}})={\mathcal{W}}({{\mbox{$\displaystyle\alpha$}}})-2.$
By construction, the $\Delta$-orbit of ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{$\displaystyle\alpha$}}$ consists of pairwise disjoint rectangles. Therefore, ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{$\displaystyle\alpha$}}$ projects injectively onto a rectangle ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha}$ in $V$. Moreover, different ${\boldsymbol{\Pi}}^{\operatorname{can}}_{{\mbox{$\displaystyle\alpha$}}_{i}}$ project to pairwise disjoint ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha_{i}}$.
If ${\mathcal{W}}({{\mbox{$\displaystyle\alpha$}}})\leq 2$ we let ${\mathcal{W}}^{\operatorname{can}}(\alpha)=0$ and ${\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)\coloneqq\emptyset$. Arcs with positive weight form the canonical weighted arc diagram (WAD) on $V$:
@@ 3.3.5. Local WADs and the thick-thin decomposition
See [L3, §7.6] for a reference of the following fact.
@@ 3.3.8. WAD under covering
Assume that $f\colon U\to V$ as a finite-degree covering between surfaces with boundaries, where $\chi(V)<0$. Then, see [K, Lemma 3.3]
@@ 3.4.2. Definitions
A pseudo-quadratic-like map ($\psi^{\bullet}$-ql map”) is a pair of holomorphic maps
@@
    $f\colon U\to V$ is a double branched covering (we usually normalize it so that its critical point is located at $0$);
@@ 3.4.3. $\psi^{\bullet}$-ql renormalization
$U_{n,i}\coloneqq{\mathbb{A}}\left(U\setminus f_{n,i}^{-1}({\boldsymbol{\mathfrak{B}}}^{[n]}),\ \partial{\mathfrak{B}}^{[n](1)}_{i}\right)\bigcup_{\partial{\mathfrak{B}}^{[n](1)}_{i}}{\mathfrak{B}}^{[n](1)}_{i}$
@@
We will write suppress index $0$” for $F_{n,0}$:
@@
$\Upsilon^{\prime}=\iota^{-1}(\Upsilon)\simeq\Upsilon=\Upsilon(f_{X}),\hskip 42.67912pt\text{ i.e., }\Upsilon\text{ is $\displaystyle\iota$-proper}.$
For example, one can take $\Upsilon={\mathfrak{T}}$ or $\Upsilon={\mathfrak{T}}\cup{\boldsymbol{\mathfrak{B}}}^{[m]}$. To define $\psi^{\Upsilon}$-renormalization”, one should extend a ql prerenormalization $f_{n,i}\colon{X_{n,i}\to Y_{n,i}}$ along all curves in $V\setminus\bigcup_{j}\Upsilon_{f_{n,j}}$. Moreover, $\Upsilon_{f_{n,j}}$ can often be enlarged, see §3.4.5.
@@ 3.4.4. Sup-Chain Rule for renormalization domains
[…]
@@ 3.4.5. Unbalanced renormalization from $\psi^{\bullet}$ to $\psi$-ql maps
We recall from §2.4.5 that the set ${\boldsymbol{\Upsilon}}$ is inhomogeneous: $f_{X}$ does not permute components of ${\boldsymbol{\Upsilon}}$ .
@@ 3.4.7. Sup-Chain rule for iterations
(3.20) $\rho^{\prime}\colon V\setminus{\mathfrak{K}}_{F}\hookrightarrow M\setminus{\mathfrak{K}}_{G}\overset{\rho}{\longrightarrow}B\setminus{\boldsymbol{\Upsilon}}$
@@ 3.5.2. WADs of $F$
Assume that ${\boldsymbol{\mathfrak{B}}}^{[n]}$ is well-defined, i.e., $F$ is $n+1$ renormalizable. Let us consider the canonical WAD of $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$, $k\geq m$:
${\mathcal{A}}^{(m),k}\equiv{\mathcal{A}}^{[n],(m),k}\coloneqq{\operatorname{WAD}}\left(U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\right).$
@@
  • ${\mathcal{A}}_{i}^{(m),k}$ to be the local WAD for ${\mathfrak{B}}^{[n],(m),k}_{i}$, see §3.3.5: ${\mathcal{A}}_{i}^{(m),k}$ consisting of arcs adjacent to ${\mathfrak{B}}^{[n],(m),k}_{i}$ such that the weights of self-arcs adjacent to ${\mathfrak{B}}^{[n],(m),k}_{i}$ are doubled;
@@
Assume that in the dynamical plane of $F={\mathcal{R}}^{n_{0}\bullet}G\colon U\rightrightarrows V$, there is a non-trivial horizontal lamination of curves ${\mathcal{L}}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ emerging from ${\mathfrak{B}}^{[n]}_{0}$ and landing at ${\boldsymbol{\mathfrak{B}}}^{[n]}$. Then
@@ 3.5.3. WADs and ${\mathcal{W}}_{\bullet}(F_{n,i})$
(3.24) $\frac{1}{2}{\mathcal{W}}_{\bullet}(F_{n,0})\leq{\mathcal{W}}_{\bullet}(F_{n,i})\leq{\mathcal{W}}_{\bullet}(F_{n,0})\hskip 19.91684pt\text{ for all }\ i.$
Indeed, let $\gamma_{i}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ be the peripheral hyperbolic geodesic around ${\mathfrak{B}}^{[n]}_{i}$. Then the hyperoblic length of $|\gamma_{i}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$ is proportional to ${\mathcal{W}}_{\bullet}(F_{n,i})$. Let $\gamma^{1}_{i}$ be the lift of $\gamma_{i}$ under $f\colon U^{1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}\to V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ so that $\gamma^{1}_{i}$ is the peripheral hyperbolic geodesic around ${\mathfrak{B}}^{[n],(1)}_{i-1}$, where the subscript is $\operatorname{mod}p_{n}$. Counting the degree of $f\colon\gamma^{1}_{i}\to\gamma_{i}$, we obtain
  • $|\gamma^{1}_{i}|_{U^{1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}=|\gamma_{i}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$ if $i\not=1$; and
  • $\frac{1}{2}|\gamma^{1}_{1}|_{U^{1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}=|\gamma_{1}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$.
Finally, $|\gamma^{1}_{i}|_{U^{1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}\geq|\gamma_{i-1}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$; this implies (3.24).
@@
Assume that $m\leq k$. By Lemma 3.6, the set ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\subset U^{k}$ is $\iota$-proper. Let $\gamma$ be a proper curve in $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ emerging from ${\mathfrak{B}}^{[n],(m)}_{a}$. Applying $\iota^{-s}$ along $\gamma$, we construct, see (3.26) below, its restriction $(\iota^{*})^{s}[\gamma]$ which is a proper curve in $U^{k+s}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$. Moreover,
@@
The curve $\gamma\in{\mathcal{F}}^{\operatorname{full}}(R)$ becomes vertical under the restriction by $(\iota^{s})^{*}$ if and only if $\gamma(t)$ hits $\iota^{s}\big(\partial U^{k+s}\big)$. This is impossible because $\gamma_{1}(t),\gamma_{2}(t)$ do not hit $\iota^{s}\big(\partial U^{k+s}\big)$. ∎
@@ 3.5.6. Monotonicity of the ${\mathcal{A}}^{k}_{\mathrm{hor}}$
It follows essentially from Lemma 3.3 (by lifting $\iota$ to the universal coverings, see [K, §3.5] for reference) that for $k\geq m$
(3.27) ${\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}}\geq{\mathcal{A}}^{[n],(m),k+1}_{\mathrm{hor}}\hskip 14.22636pt\text{ and }\hskip 14.22636pt{\mathcal{A}}^{[n],(m),k}_{i,{\mathrm{hor}}}\geq{\mathcal{A}}_{i,{\mathrm{hor}}}^{[n],(m),k+1}.$
@@
It follows from Decomposition (3.28) that there is a combinatorial constant $C\equiv C_{p_{n},m}$ (essentially, the maximal number of possible arcs rel ${\boldsymbol{\mathfrak{B}}}^{[n],(m+1)}$) such that for all $M>1$
(3.29) ${\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}}-MC{\ \ }{\ \ }\text{ is \emph{dominated} by }{\ \ }{\ \ }{\mathcal{A}}^{[n],(m+1),k}_{\mathrm{hor}}-M$
in the following sense, see §3.3.6 for notation. For every arc $\alpha$ in ${\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}}-MC$, there is a vertical sublamination
${\mathcal{L}}_{\alpha}\subset{\mathcal{F}}(e)\equiv{\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha){\ \ }{\ \ }\text{ with }{\ \ }{\ \ }{\mathcal{W}}({\mathcal{L}}_{\alpha})\geq{\mathcal{W}}(\alpha)-MC$
@@
(3.30) $\gamma_{i}\in{\mathcal{F}}^{\operatorname{full}}\left({\mathcal{A}}^{[n],(m+1),k}_{\mathrm{hor}}-M\right).$
The lamination ${\mathcal{L}}_{a}$ is constructed by removing all $\gamma$ from ${\mathcal{F}}(\alpha)$ that do not satisfy (3.30) – the width of removed curves is bounded by $MC$.